OpenStudy (anonymous):

I have quadratic equation 5x^2-20x=0, i've factored out the 5x and now have 5x(x-4)=0, but don't know how to get the two solutions from that. Help??? I have quadratic equation 5x^2-20x=0, i've factored out the 5x and now have 5x(x-4)=0, but don't know how to get the two solutions from that. Help??? @Mathematics

OpenStudy (anonymous):

x = 0 x = 4 5(0) ( x-4) = 0 5(x) ( (4) -(4)) = 0

OpenStudy (saifoo.khan):

5x(x-4) = 0

OpenStudy (lilg132):

5x = 0 and x - 4 = 0 which is x = 4 u have two answers x = 0 and x = 4

OpenStudy (anonymous):

thank you! all were helpful!